Use matrix techniques to solve the initial value problem -6 2. X X X(0) = 18 -3 1
Answer #1

x=1-62 og [3]* *(0) = [1] let A={:} & first all we will find Eigenvalue and Eigen vector for A. = (-d-6)(1-0) +620 = 2² +50=